# Scaling and Load-Balancing Equi-Joins

### v2

@@ Abstract
This paper proposes Adaptive-Multistage-Join (AM-Join) for scalable and fast equi-joins in distributed shared-nothing architectures. AM-Join utilizes (a) Tree-Join, a proposed novel algorithm that scales well when the joined tables share hot keys, and (b) Broadcast-Join, the fastest known algorithm when joining keys that are hot in only one table.
@@
1 Introduction
@@
Retrieval of information from two database tables is critical for data processing, and impacts the computational cost and the response time of queries. In the most general case, this operation entails carrying out a cross join of the two tables. The more common case is computing an equi-join, where two records in the two tables are joined if and only if equality holds between their keys (values in the join column(s)). The algorithms for equi-joins have been optimized regularly since the inception of the database community [31, 32, 78, 16, 28, 36, 40, 76, 53, 79, 39].
@@
Significant research has been done to enhance the sequential equi-join algorithms on multi-core processors [51, 18, 4, 10, 9, 77, 13, 11] and on GPUs [46, 50, 22, 44, 80, 66, 73, 67]. However, the proliferation of data collection and analysis poses a challenge to sequential join algorithms that are limited by the number of threads supported by the processing units. Scaling equi-joins had to progress through distributed architectures, which is the direction adopted in this work.
@@
This work is motivated by equi-joining industry-scale skewed datasets in a novel way. We tackle natural self-joins[31, 32] at the intersection of equi-joins [33], inner-joins, and self-joins [48]. This join semantic is an integral operation in the similarity-based join algorithms (e.g., [63]) used for fraud detection. While natural self-joins can be more efficient to perform than the general equi-joins, the state-of-the-art equi-join algorithms failed to scale to our large and skewed datasets. This motivated us to develop the fast, efficient, and scalable Adaptive-Multistage-Join (AM-Join) that scales, not only to our specialized use case, but also to the most challenging equi-joins.
@@
We first propose Tree-Join, a novel algorithm that scales well by distributing the load of joining a key that is hot (i.e., high-frequency or shared by a large number of records) in both tables to multiple executors. Such keys are the scalability bottleneck of most of the state-of-the-art distributed algorithms. We give special attention to balancing the load among the executors throughout the Tree-Join execution.
@@
We then tackle Small-Large joins, where one table fits in memory and the other can be up to orders of magnitude larger. We devise the Index-Broadcast-Join (IB-Join) family for Small-Large joins, and show analytically their outer-join variants improve on the state-of-the-art Small-Large outer-join algorithms [91, 27].
@@
The Tree-Join, and the Broadcast-Join algorithms are the building blocks of AM-Join. AM-Join achieves (a) high scalability by utilizing Tree-Join that distributes the load of joining a key that is hot in both relations to multiple executors, and (b) fast execution by utilizing the Broadcast-Join algorithms that reduce the network load when joining keys that are hot in only one relation. AM-Join extends to all outer-joins elegantly without record deduplication or custom table partitioning, unlike the state-of-the-art industry-scale algorithms [21]. The outer-join variants of AM-Join achieves fast execution by utilizing the IB-Join family of algorithms.
@@
All the proposed algorithms use the basic MapReduce primitives only, and hence can be adopted on any shared-nothing architecture. We implemented a MapReduce version using Spark [95]. Our evaluation highlights the improved performance and scalability of AM-Join when applied to the general equi-joins. When compared to the sate-of-the-art algorithms [21, 41], AM-Join executed comparably fast on weakly-skewed synthetic tables and can join more-skewed or orders-of-magnitude bigger tables, including our real-data tables. These advantages are even more pronounced when applying the join algorithms to natural self-joins. The proposed IB-Join outer-join algorithm executed much faster than the algorithms in [91, 27].
@@
This manuscript is an extended version of a prior paper [60].
2 Formalization
@@
$Key$ The join attribute(s)/column(s).
@@
$\ell$ $\ell_{\mathcal{R}}$, $\ell_{\mathcal{S}}$ The number of records, a.k.a., frequency,
@@
Table 1: The symbols used in the paper.
2.1 Equi-Joins, Inner-Joins and Self-Joins
Figure 1: An example for joining two relations. (a) shows the input relations. (b) through (e) show the results of the inner, left-outer, right-outer, and full-outer-joins, respectively. The tables are sorted for readability, but sorting is not guaranteed in practice.
The equi-join operation combines columns from two tables, a.k.a. relations in the relational database community, based on the equality of the column(s), a.k.a., join attribute(s). We focus on the case where two relations are joined. Other work, e.g., [2, 64, 29, 1], focused on extending equi-joins to multiple relations.
@@ 2.1 Equi-Joins, Inner-Joins and Self-Joins
In case $\mathcal{R}$ and $\mathcal{S}$ are the same table, this is called self-join [48]. A table can be self-joined using different join attributes, $Key_{1}$ and $Key_{2}$ on the two sides of the join, captured by the notation $\mathcal{R}\Join_{key_{1_{r_{1}}}\theta key_{2_{r_{2}}}}\mathcal{R}$, where $\theta$ is the function that has to hold between the $Key_{1}$ and $Key_{2}$ attribute(s) of records $r_{1}$ and $r_{2}$ respectively for the pair $r_{1}$-$r_{2}$ to belong to $\mathcal{Q}$. If the same join attributes are used on both sides of the self-join, i.e., $Key1=Key2=Key$, and $\theta$ is the equality function, i.e., the join condition is $key_{r_{1}}=key_{r_{2}}$, then this is an inner-join by definition, and is captured by the notation $\mathcal{R}\Join_{key_{r_{1}}=key_{r_{2}}}\mathcal{R}$, or $\mathcal{R}\Join_{Key}\mathcal{R}$ for short. We call this special case natural self-join. The $\mathcal{Q}$ produced by a natural self-join contains duplicate pairs of records, but in reverse order. To eliminate the redundancy in $\mathcal{Q}$, we drop the join completeness: (a) a pair $r-r$ should be in $\mathcal{Q}$ exactly once, and (b) a pair $r_{2}$-$r_{1}$ should not be in $\mathcal{Q}$, if $r_{1}\neq r_{2}$ and $r_{1}$-$r_{2}$ is in $\mathcal{Q}$.
2.2 Popular Distributed-Processing Frameworks
MapReduce [37] is a popular framework with built-in fault tolerance. It allows developers, with minimal effort, to scale data-processing algorithms to shared-nothing clusters, as long as the algorithms can be expressed in terms of the functional programing primitives mapRec and reduceRecs.
@@ 2.2 Popular Distributed-Processing Frameworks
In addition to $key_{2}$, the mapper can optionally output tuples by a secondary key. The reduce_value_list would also be sorted by the secondary key in that case [^1] Secondary keys are not supported by the public version of MapReduce, Hadoop [6]. Two ways to support secondary keys were proposed in [58]. The first entails loading the entire reduce_value_list in the reducer memory, and the second entails rewriting the shuffler. The second solution is more scalable but incurs higher engineering cost.. Partial reducing can happen at the mappers, which is known as combining to reduce the network load. For more flexibility, the MapReduce framework allows for loading external data when mapping or reducing. However, to preserve the determinism and the purity of the mapRec and reduceRecs functions, loading is restricted to the beginning of each operation.
The Spark framework [95] is currently the de facto industry standard for distributing data processing on shared-nothing clusters. Spark offers the functionality of MapReduce [^2] While Spark uses the same nomenclature of MapReduce, the MapReduce map function is called flatMap in Spark. We use the MapReduce notation, introduced in [37]., as well as convenience utilities that are not part of MapReduce, but can be built using MapReduce. One example is performing tree-like aggregation of all the records in a dataset [54]. The treeAggregate operation can be implemented using a series of MapReduce stages, where each stage aggregates the records in a set of data partitions, and the aggregates are then aggregated further in the next stage, and so on, until the final aggregate is collected at the driver.
@@
Spark is designed as an in-memory framework. It relies on the executor memory, rather than their local disks, to store and process data. This in-memory processing is a key feature that gives Spark its speed and efficiency, but constraints the size of the records (including the reduce_value_lists) that can be input or output by any executor at any time. Even more, while Spark can persist dataset partitions on local disks or checkpoint them on a distributed file system, when Spark processes a data partition, it loads the entire partition in memory [^3] The reader is referred to the Spark programming guide, https://spark.apache.org/docs/3.0.2/rdd-programming-guide.html#rdd-persistence..
3 Related Work
@@
We review the distributed algorithms that are generally applicable to equi-joins, and those targeting Small-Large outer-joins. We only review equi-joins in homogenous shared-nothing systems, where the processing powers of the executors are comparable. We neither review the algorithms devised for heterogenous architectures (e.g., [85]) nor in streaming systems (e.g., [35, 43, 57]), nor the approximate equi-join algorithms (e.g., [70]).
3.1 General Distributed Equi-Joins
@@
The work most relevant to ours is that of distributing equi-joins on MapReduce [37] and on general shared-nothing [82] architectures.
3.1.1 MapReduce Equi-Joins
@@
The seminal work in [19] explained Map-Side join (a.k.a., Broadcast or Duplication join in the distributed-architecture community), which is discussed in more details in § 5. This work also extended the Sort-Merge join [20] to the MapReduce framework. It also proposed using secondary keys to alleviate the bottleneck of loading the records from both joined relations in memory. This Sort-Merge join is considered a Reduce-Side join (a.k.a., Shuffle join in the distributed-architecture community). However, the work in [19] overlooks the skew in key popularity.
@@
The SAND Join [8], a rediscovery of [89], extends the Reduce-Side joins by partitioning data across reducers based on a sample. The work in [65] improves this key-range partitioning using quantile computation. It extends key-range partitioning to combinations of keys in the join results, and maps combinations of keys to specific reducers to achieve load-balancing. The algorithms in [87, 41] use sampling to build the cross–relation histogram, and balance the load between the executors based on the sizes of the relations and the estimated size of the join results.
@@
There are major drawbacks with key-range-division approaches [89, 8, 65, 87, 41]. In the pre-join step, the key-range computation and communication over the network to the mappers incurs significant computational [24] and communication cost [94]. The computational and network bottlenecks are more pronounced when computing key-ranges for combinations of attributes [65] [^4] This problem was the focus of [5] in the context of estimating join results size [83, 88, 86, 23].. These approaches implicitly assume that the keys are distributed evenly within each key range, which is rarely true for highly-diverse key spaces, highly-skewed data, or practical key-ranges.
@@
The multi-executor-per-key Shuffle-Join algorithms in [2, 15] splits the processing of each hot key among multiple executors. The executors that process each key are organized in a grid whose dimensions, length and width, are proportional to the frequencies of the key on $\mathcal{R}$ and $\mathcal{S}$, respectively. Each record from $\mathcal{R}$ is sent to all the executors of a random grid row, and each record from $\mathcal{S}$ is sent to all the executors of a random grid column. Each executor then outputs the cartesian product of all pairs it receives.
@@
Figure 2: A schematic example for joining a specific hot key, $k_{1}$, using the multi-executor-per-key Shuffle-Join algorithms [2, 65, 15, 56]. (a) The $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. (b) The records with key $k_{1}$ replicated among the rows and columns of the grid of executors assigned to $k_{1}$.
@@
This gird approach was modified into a triangle for natural self-joins in the context of record deduplication [30] to reduce roughly half the processing cost. If Fig. 2 was to illustrate a self-join, both $\mathcal{R}$ and $\mathcal{S}$ in Fig. 3(a) would be the same relation. The executor grid in Fig. 2(b) would be an upper triangle matrix, and the lower triangle cells would not be assigned to executors.
ExpVar-Join [56] goes a step further by trying to find the optimal grid dimensions (or triangle side length for natural self-joins) for each hot key that would balance the load on all executors. The algorithm executes a loop, and in each iteration it computes the expected runtime and variance in the executor load, and hence the name of the algorithm. The search loop is executed until the expected runtime in the last $n$ iterations stabilizes (varies by less than $1\%$), where $n$ is the number of executors. In each iteration, the algorithm finds the hot key dimension that would reduce the load variance the most if assigned an extra executor. The estimation is done using a greedy approach that has near-optimality guarantees assuming (a) a custom shuffle is implemented, and (b) each grid dimension for each key can grow until it is within a factor of $O(\lambda)$ of the key’s frequency, where $\lambda$ is the relative speed of the network vs. a local disk. The runtime is estimated by building a linear regression model that needs to be executed in a pre-processing phase on a sample of the data.
@@
3.1.2 Shared-Nothing Equi-Joins

### v3

@@ Abstract.
This paper, an extended version of (Metwally, 2022), proposes Adaptive-Multistage-Join (AM-Join) for scalable and fast equi-joins in distributed shared-nothing architectures. AM-Join utilizes (a) Tree-Join, a proposed novel algorithm that scales well when the joined tables share hot keys, and (b) Broadcast-Join, the fastest known algorithm when joining keys that are hot in only one table.
@@
Big Data; Data Skew; Distributed Algorithms; Load Balancing
copyright: acmlicensedjournal: TODSjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3722102ccs: Information systems Join algorithmsccs: Theory of computation MapReduce algorithms
1. Introduction
@@
Retrieval of information from two database tables is critical for data processing, and impacts the computational cost and the response time of queries. In the most general case, this operation entails carrying out a cross join of the two tables. The more common case is computing an equi-join, where two records in the two tables are joined if and only if equality holds between their keys (values in the join column(s)). The algorithms for equi-joins have been optimized regularly since the inception of the database community (Codd, 1969, 1970; Selinger et al., 1979; Bernstein et al., 1981; Cheung, 1982; Date, 1983; DeWitt et al., 1987; Schneider and DeWitt, 1989; Lakshmi and Yu, 1990; Shasha and Wang, 1991; DeWitt et al., 1992).
@@
Significant research has been done to enhance the sequential equi-join algorithms on multi-core processors (Kim et al., 2009; Blanas et al., 2011; Albutiu et al., 2012; Balkesen et al., 2013b, a; Schuh et al., 2016; Barthels et al., 2015; Bandle et al., 2021) and on GPUs (He et al., 2008; Kaldewey et al., 2012; Chen and Prasanna, 2016; Guo et al., 2019; Sioulas et al., 2019; Paul et al., 2020; Rui et al., 2020; Paul et al., 2021). However, the proliferation of data collection and analysis poses a challenge to sequential join algorithms that are limited by the number of threads supported by the processing units. Scaling equi-joins had to progress through distributed architectures, which is the direction adopted in this work.
@@
This work is motivated by equi-joining industry-scale skewed datasets in a novel way. We tackle natural self-joins (Codd, 1969, 1970) at the intersection of equi-joins (Codd, 1972), inner-joins, and self-joins (Hursch, 1989). This join semantic is an integral operation in the similarity-based join algorithms (e.g., (Metwally and Faloutsos, 2012; Metwally and Shum, 2024)) used for fraud detection. While natural self-joins can be more efficient to perform than the general equi-joins, the state-of-the-art equi-join algorithms failed to scale to our large and skewed datasets. This motivated us to develop the fast, efficient, and scalable Adaptive-Multistage-Join (AM-Join) that scales, not only to our specialized use case, but also to the most challenging equi-joins.
@@
In this paper, an extended version of (Metwally, 2022), we first propose Tree-Join, a novel algorithm that scales well by distributing the load of joining a key that is hot (i.e., high-frequency or shared by a large number of records) in both tables to multiple executors. Such keys are the scalability bottleneck of most of the state-of-the-art distributed algorithms. We give special attention to balancing the load among the executors throughout the Tree-Join execution.
@@
We then tackle Small-Large joins, where one table fits in memory and the other can be up to orders of magnitude larger. We devise the Index-Broadcast-Join (IB-Join) family for Small-Large joins, and show analytically their outer-join variants improve on the state-of-the-art Small-Large outer-join algorithms (Xu and Kostamaa, 2010; Cheng et al., 2017).
@@
The Tree-Join, and the Broadcast-Join algorithms are the building blocks of AM-Join. AM-Join achieves (a) high scalability by utilizing Tree-Join that distributes the load of joining a key that is hot in both relations to multiple executors, and (b) fast execution by utilizing the Broadcast-Join algorithms that reduce the network load when joining keys that are hot in only one relation. AM-Join extends to all outer-joins elegantly without record deduplication or custom table partitioning, unlike the state-of-the-art industry-scale algorithms (Bruno et al., 2014). The outer-join variants of AM-Join achieves fast execution by utilizing the IB-Join family of algorithms.
@@
All the proposed algorithms use the basic MapReduce primitives only, and hence can be adopted on any shared-nothing architecture. We implemented a MapReduce version using Spark (Zaharia et al., 2010). Our evaluation highlights the improved performance and scalability of AM-Join when applied to the general equi-joins. When compared to the sate-of-the-art algorithms (Bruno et al., 2014; Gavagsaz et al., 2019), AM-Join executed comparably fast on weakly-skewed synthetic tables and can join more-skewed or orders-of-magnitude bigger tables, including our real-data tables. These advantages are even more pronounced when applying the join algorithms to natural self-joins. The proposed IB-Join outer-join algorithm executed much faster than the algorithms in (Xu and Kostamaa, 2010; Cheng et al., 2017).
@@
2. Formalization
@@
$Key$. The join attribute(s)/column(s).
@@
$\ell$, $\ell_{\mathcal{R}}$, $\ell_{\mathcal{S}}$ The number of records, a.k.a., frequency,
@@
2.1. Equi-Joins, Inner-Joins and Self-Joins
Figure 1. An example for joining two relations. (a) shows the input relations. (b) through (e) show the results of the inner, left-outer, right-outer, and full-outer-joins, respectively. The tables are sorted for readability, but sorting is not guaranteed in practice.
The equi-join operation combines columns from two tables, a.k.a. relations in the relational database community, based on the equality of the column(s), a.k.a., join attribute(s). We focus on the case where two relations are joined. Other work, e.g., (Afrati and Ullman, 2010; Nica et al., 2012; Chu et al., 2015; Afrati et al., 2018), focused on extending equi-joins to multiple relations.
@@ 2.1. Equi-Joins, Inner-Joins and Self-Joins
In case $\mathcal{R}$ and $\mathcal{S}$ are the same table, this is called self-join (Hursch, 1989). A table can be self-joined using different join attributes, $Key_{1}$ and $Key_{2}$ on the two sides of the join, captured by the notation $\mathcal{R}\Join_{key_{1_{r_{1}}}\theta key_{2_{r_{2}}}}\mathcal{R}$, where $\theta$ is the function that has to hold between the $Key_{1}$ and $Key_{2}$ attribute(s) of records $r_{1}$ and $r_{2}$ respectively for the pair $r_{1}$-$r_{2}$ to belong to $\mathcal{Q}$. If the same join attributes are used on both sides of the self-join, i.e., $Key1=Key2=Key$, and $\theta$ is the equality function, i.e., the join condition is $key_{r_{1}}=key_{r_{2}}$, then this is an inner-join by definition, and is captured by the notation $\mathcal{R}\Join_{key_{r_{1}}=key_{r_{2}}}\mathcal{R}$, or $\mathcal{R}\Join_{Key}\mathcal{R}$ for short. We call this special case natural self-join. The $\mathcal{Q}$ produced by a natural self-join contains duplicate pairs of records, but in reverse order. To eliminate the redundancy in $\mathcal{Q}$, we drop the join completeness: (a) a pair $r-r$ should be in $\mathcal{Q}$ exactly once, and (b) a pair $r_{2}$-$r_{1}$ should not be in $\mathcal{Q}$, if $r_{1}\neq r_{2}$ and $r_{1}$-$r_{2}$ is in $\mathcal{Q}$.
2.2. Popular Distributed-Processing Frameworks
MapReduce (Dean and Ghemawat, 2008) is a popular framework with built-in fault tolerance. It allows developers, with minimal effort, to scale data-processing algorithms to shared-nothing clusters, as long as the algorithms can be expressed in terms of the functional programing primitives mapRec and reduceRecs.
@@ 2.2. Popular Distributed-Processing Frameworks
In addition to $key_{2}$, the mapper can optionally output tuples by a secondary key. The reduce_value_list would also be sorted by the secondary key in that case [^1] Secondary keys are not supported by the public version of MapReduce, Hadoop (Apache Hadoop, [n. d.]). Two ways to support secondary keys were proposed in (Lin and Dyer, 2010). The first entails loading the entire reduce_value_list in the reducer memory, and the second entails rewriting the shuffler. The second solution is more scalable but incurs higher engineering cost.. Partial reducing can happen at the mappers, which is known as combining to reduce the network load. For more flexibility, the MapReduce framework allows for loading external data when mapping or reducing. However, to preserve the determinism and the purity of the mapRec and reduceRecs functions, loading is restricted to the beginning of each operation. The Spark framework (Zaharia et al., 2010) is currently the de facto industry standard for distributing data processing on shared-nothing clusters. Spark offers the functionality of MapReduce [^2] While Spark uses the same nomenclature of MapReduce, the MapReduce map function is called flatMap in Spark. We use the MapReduce notation, introduced in (Dean and Ghemawat, 2008)., as well as convenience utilities that are not part of MapReduce, but can be built using MapReduce. One example is performing tree-like aggregation of all the records in a dataset (Lämmel, 2008). The treeAggregate operation can be implemented using a series of MapReduce stages, where each stage aggregates the records in a set of data partitions, and the aggregates are then aggregated further in the next stage, and so on, until the final aggregate is collected at the driver.
@@
Spark is designed as an in-memory framework. It relies on the executor memory, rather than their local disks, to store and process data. This in-memory processing is a key feature that gives Spark its speed and efficiency, but constrains the size of the records (including the reduce_value_lists) that can be input or output by any executor at any time. Even more, while Spark can persist dataset partitions on local disks or checkpoint them on a distributed file system, when Spark processes a data partition, it loads the entire partition in memory [^3] The reader is referred to the Spark programming guide, https://spark.apache.org/docs/latest/rdd-programming-guide.html#rdd-persistence..
3. Related Work
@@
We review the distributed algorithms that are generally applicable to equi-joins, and those targeting Small-Large outer-joins. We only review equi-joins in homogenous shared-nothing systems, where the processing powers of the executors are comparable. We neither review the algorithms devised for heterogeneous architectures (e.g., (Tian et al., 2016)) nor in streaming systems (e.g., (Das et al., 2003; Gulisano et al., 2016; Lin et al., 2015)), nor the approximate equi-join algorithms (e.g., (Quoc et al., 2018)).
3.1. General Distributed Equi-Joins
@@
The work most relevant to ours is that of distributing equi-joins on MapReduce (Dean and Ghemawat, 2008) and on general shared-nothing (Stonebraker, 1986) architectures.
3.1.1. MapReduce Equi-Joins
@@
The seminal work in (Blanas et al., 2010) explained Map-Side join (a.k.a., Broadcast or Duplication join in the distributed-architecture community), which is discussed in more detail in § 5. This work also extended the Sort-Merge join (Blasgen and Eswaran, 1977) to the MapReduce framework. It also proposed using secondary keys to alleviate the bottleneck of loading the records from both joined relations in memory. This Sort-Merge join is considered a Reduce-Side join (a.k.a., Shuffle join in the distributed-architecture community). However, the work in (Blanas et al., 2010) overlooks the skew in key popularity.
@@
The SAND Join (Atta et al., 2011), a rediscovery of (Wolf et al., 1993), extends the Reduce-Side joins by partitioning data across reducers based on a sample. The work in (Okcan and Riedewald, 2011) improves this key-range partitioning using quantile computation. It extends key-range partitioning to combinations of keys in the join results, and maps combinations of keys to specific reducers to achieve load-balancing. The algorithms in (Vitorovic et al., 2016; Gavagsaz et al., 2019) use sampling to build the cross–relation histogram, and balance the load between the executors based on the sizes of the relations and the estimated size of the join results.
@@
There are major drawbacks with key-range-division approaches (Wolf et al., 1993; Atta et al., 2011; Okcan and Riedewald, 2011; Vitorovic et al., 2016; Gavagsaz et al., 2019). In the pre-join step, the key-range computation and communication over the network to the mappers incurs significant computational (Chen and Zhang, 2020) and communication cost (Yi and Zhang, 2013). The computational and network bottlenecks are more pronounced when computing key-ranges for combinations of attributes (Okcan and Riedewald, 2011) [^4] This problem was the focus of (Alway and Nica, 2016) in the context of estimating join results size (Sun et al., 1993; Wang et al., 2003; Vengerov et al., 2015; Chen and Yi, 2017).. These approaches implicitly assume that the keys are distributed evenly within each key range, which is rarely true for highly-diverse key spaces, highly-skewed data, or practical key-ranges.
@@
The multi-executor-per-key Shuffle-Join algorithms in (Afrati and Ullman, 2010; Beame et al., 2014) splits the processing of each hot key among multiple executors. The executors that process each key are organized in a grid whose dimensions, length and width, are proportional to the frequencies of the key on $\mathcal{R}$ and $\mathcal{S}$, respectively. Each record from $\mathcal{R}$ is sent to all the executors of a random grid row, and each record from $\mathcal{S}$ is sent to all the executors of a random grid column. Each executor then outputs the cartesian product of all pairs it receives.
@@
Figure 2. A schematic example for joining a specific hot key, $k_{1}$, using the multi-executor-per-key Shuffle-Join algorithms (Afrati and Ullman, 2010; Okcan and Riedewald, 2011; Beame et al., 2014; Li et al., 2018). (a) The $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. (b) The records with key $k_{1}$ replicated among the rows and columns of the grid of executors assigned to $k_{1}$.
@@
This gird approach was modified into a triangle for natural self-joins in the context of record deduplication (Chu et al., 2016) to reduce roughly half the processing cost. If Fig. 2 was to illustrate a self-join, both $\mathcal{R}$ and $\mathcal{S}$ in Fig. 3(a) would be the same relation. The executor grid in Fig. 2(b) would be an upper triangular matrix, and the lower triangle cells would not be assigned to executors.
ExpVar-Join (Li et al., 2018) goes a step further by trying to find the optimal grid dimensions (or triangle side length for natural self-joins) for each hot key that would balance the load on all executors. The algorithm executes a loop, and in each iteration it computes the expected runtime and variance in the executor load, and hence the name of the algorithm. The search loop is executed until the expected runtime in the last $n$ iterations stabilizes (varies by less than $1\%$), where $n$ is the number of executors. In each iteration, the algorithm finds the hot key dimension that would reduce the load variance the most if assigned an extra executor. The estimation is done using a greedy approach that has near-optimality guarantees assuming (a) a custom shuffle is implemented, and (b) each grid dimension for each key can grow until it is within a factor of $O(\lambda)$ of the key’s frequency, where $\lambda$ is the relative speed of the network vs. a local disk. The runtime is estimated by building a linear regression model that needs to be executed in a pre-processing phase on a sample of the data.
@@
3.1.2. Shared-Nothing Equi-Joins